You are asked to evaluate the following two projects for the Norton corporation. Use a discount rate of 10 percent. Use Appendix B for an approximate answer but calculate your final answer using the formula and financial calculator methods.

  

Project X (Videotapes
of the Weather Report)
($46,000 Investment)
Project Y (Slow-Motion
Replays of Commercials)
($80,000 Investment)
Year Cash Flow Year Cash Flow   
1 $ 30,000 1 $ 40,000
2 28,000 2 33,000
3 20,000 3 34,000
4 18,600 4 36,000

  

a.

Calculate the profitability index for project X. (Do not round intermediate calculations and round your answer to 2 decimal places.)

    

  Profitability index   

  

b.

Calculate the profitability index for project Y. (Do not round intermediate calculations and round your answer to 2 decimal places.)

  

  Profitability index   

  

c.

Which project would you select?

  
Project X
Project Y

Hints

References

eBook & Resources
Answer #1

Profitability index= 1+NPV/Initial Inv

PROJECT X
YEAR CASH FLOWs DCF@10% DISC Cash Flows
0 -46000 1 -46000.00
1 30000 0.91 27272.73
2 28000 0.83 23140.50
3 20000 0.75 15026.30
4 18600 0.68 12704.05
NPV 32143.57

Profitability index=1+32143.57/46000=1.69

PROJECT Y
YEAR CASH FLOWs DCF@10% DISC Cash Flows
0 -80000 1 -80000.00
1 40000 0.91 36363.64
2 33000 0.83 27272.73
3 34000 0.75 25544.70
4 36000 0.68 24588.48
NPV 33769.55

profitability index= 1+33769.55/80000=1.42

Accept a project if the profitability index is greater than 1, stay indifferent if the profitability index is zero and don't accept a project if the profitability index is below 1.

since both th projects have above 1 profitability index decision must be taken with the help of npv. project with greter npv should be accepted.

project Y having greater npv than project X

PROJECT Y SHOULD BE SELECTED